OpenStudy (anonymous):

which of the following points are on the line given by the equation y=4x+2?

geerky42 (geerky42):

One of the ways to solve it is plugging in x value then solve for y and see if it matches

OpenStudy (anonymous):

well my options are a. -2,-6 b. 1,6 c. -2,-10 e. 3,14

jigglypuff314 (jigglypuff314):

the options are points given in (x, y) form what geerky42 is saying is that you can plug in each option to see if it works :) for example (-2, -10) into y=4x+2 would give -10 = 4(-2) + 2 then simplify -10 = -8 + 2 -10 = -6 but -10 doesn't equal -6 so (-2, -10) would not work

jigglypuff314 (jigglypuff314):

btw are you sure that is the correct question? because then there would be 3 out of the 4 options that work o-o

OpenStudy (anonymous):

yeah it's the right one haha i tried option b so 6= 4(1) + 2 6=6 so would option b be the right answer @jigglypuff314

jigglypuff314 (jigglypuff314):

yes, it works but... but... but... if you try the other ones for some reason I think they work too :/ I'm so confused right now >,<

OpenStudy (anonymous):

oh yeah! umm idk why lmao

OpenStudy (anonymous):

oh ok it says check all that apply that makes sense now lol @jigglypuff314

jigglypuff314 (jigglypuff314):

lol okay xD

OpenStudy (anonymous):

B is right.. from the orig problem y=4x+2 the "2" is the y-intercept which is where the the line crosses the Y axis (its a positive 2) and the 4 is where it crosses the X axis so just draw a line thru those 2 numbers on each axis and then plot the points and see what falls on the line I dont have any graphing paper but if im not mistaken (1,6) should be the only point that falls on the line---- I think i did that right ;)